1.The sum of two numbers is 106, the difference of squares of two numbers is 56. Find the two numbers.
sbadaniels@hotmail.com

Answer by dfrazzetto(283)


x + y = 106
x = 106 - y
x^2 - y^2 = 56

(106 - y)^2 - y^2 = 56
11236 - 212y + y^2 - y^2 = 56
11236 - 212y = 56
212y = 11236 - 56
212y = 11180
y = 52.7358
x = 106 - y = 106 - 52.7358 = 53.2642
(53.2642)^2 - (52.7358)^2 = 2837.08 - 2781.08 = 56